nh07106:

what is 7/3 = n/21 A. 5/7 b. 3/9 c.147/3

darkknight:

7/3 = n/21 We cross multiply then 3n = 21(7) 3n = 147 So what does n equal?

nh07106:

49?

darkknight:

well 3n = 147 so n = 147/3 which is 49 yeah but 147/3 is the answer choice

nh07106:

thank ya for ur help

darkknight:

np

nh07106:

i thought i was right but i wanted ta make sure

supie:

Ok so, \[7/3=n/21\] First we must Cross multiply \[7/3=n/21\] so what you need to do is multiply \[(7)(21)=n(3)\] or 3n and 147 I think so \[147=3n\] Like always, we have to flip it around ⏎ \[3n=147\] Then divide both sides by 3(a fraction bar is a division symbol as yk) \[3n/3=147/3\]

nh07106:

ok thx

supie:

np